OpenStudy (anonymous):

okay i need help again. i;m supposed to solve this equation, i came up with two equations being equal to each other (x+2)/(x^2-16) + (x-3)/(x^2-3x-4) = (2x-3)/(x^2+5x+4)

OpenStudy (anonymous):

you made mistake somwhere, because it has one solution

OpenStudy (anonymous):

so what was my mistake?

OpenStudy (anonymous):

show your work and we will shot you mistake :P

OpenStudy (anonymous):

:) okay, i factored the denominators and got this (x+2)/(x+4)(x-4) + (x-3)/(x+1)(x-4) = (2x+3)/(x+1)(x+4) multiplying everything by the LCD which is (x+4)(x-4)(x+1) i got this (x+2)(x+1) + (x-3)(x+4) = (2x-3)(x-4) x^2+3x+2+x^2+x-12=2x^2-11x+12 2x^2+4x-10=2x^2-11x+12 that's what i have

OpenStudy (anonymous):

\[\frac{x+2}{x^2-16} + \frac{x-3}{x^2-3x-4} = \frac{2x-3}{x^2+5x+4}\] \[\frac{x+2}{(x+4)(x-4)} + \frac{x-3}{(x+1)(x-4)} = \frac{2x-3}{(x+1)(x+4)}\] \[(x+2)(x+1)+(x-3)(x+4)=(2x-3)(x-4)\] \[x^2+3x+2+x^2+x-12=2x^2-11x+12\] \[2x^2+4x-10=2x^2-11x+12\] yeah that's correct but they are not equal, subtract both sides by \(2x^2-11x+12\)

OpenStudy (anonymous):

it was my 1000th reply btw :P

OpenStudy (anonymous):

so x=22/15?

OpenStudy (anonymous):

no \[1\frac{7}{15}\] :P

OpenStudy (anonymous):

:) lol okay, thank you